# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ 4.1 Concentration-Driven Confidence Bands for Average Treatment Effects
  • The Dvoretzky–Kiefer–Wolfowitz concentration bound (which requires no tail assumptions beyond finiteness) for the order–statistic endpoints, and
  • The usual asymptotic delta‐method (CLT) for the sample means and proportions. The DKW inequality controls the uniform deviation $\sup_{x}\left\lvert F_{n}(x)-F(x)\right\rvert$ in finite samples without any distributional assumptions on $Y$ (see, e.g., Chapter 3 of Van Der Vaart et al., (1996)). This hybrid approach preserves the simplicity of the DKW envelope for the nonparametric piece while relying on asymptotic normality only for the low‐dimensional parameters.
Proposition 4 ($100(1-\alpha)\%$ hybrid confidence band under $\upalpha$–mixing).
Let $(Y_{ijt},Z_{ijt})_{i,j,t}$ be strictly stationary with $\upalpha$–mixing coefficients $\upalpha(r)$ such that
@@
$\varepsilon_{k}=(1+4C_{\upalpha})\sqrt{\frac{2\log(8/\alpha_{u})}{N_{k}}},\qquad L_{\alpha_{u}}^{(k)}=Y^{(k)}_{(1)}-\varepsilon_{k},\quad U_{\alpha_{u}}^{(k)}=Y^{(k)}_{(N_{k})}+\varepsilon_{k}.$
@@
$\nabla\mathcal{L}(\theta)=\left(p_{1},-p_{0},\delta_{1}-U_{\alpha_{u}}^{(0)},L_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top}$ (32)
$\nabla\mathcal{U}(\theta)=\left(p_{1},-p_{0},\delta_{1}-L_{\alpha_{u}}^{(0)},U_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top}$ (33)
and the covariance of the estimators $\hat{\theta}$ is given explicitly by:
Following the same logic as Corollary 1, the extension of Proposition 4 to the case of one-tail truncated latent conditional expectations entails replacing
$L_{\alpha_{u}}^{(k)}:=Y_{(1)}^{(k)}-\varepsilon_{k}\quad\text{and}\quad U^{(k)}_{\alpha_{u}}:=Y_{(N_{k})}^{(k)}+\varepsilon_{k}$
with
$L^{(k)}:=\lambda,\quad\text{and}\quad U^{(k)}_{\alpha_{u}}:=Y_{(N_{k})}^{(k)}+\tilde{\varepsilon}_{k},$
where now
$\tilde{\varepsilon}_{k}:=(1+4C_{\upalpha})\sqrt{\frac{2\log(4/\alpha_{u})}{N_{k}}},$
for $k=0,1$.
@@ 5 Monte Carlo Study
correspondingly for the Manski and Hybrid approaches. The realised outcome is
@@ 3. DGP C:
    Treatment probability:
    $\Pr(D_{it}=1\mid Y_{it}^{0})=\mathrm{logit}\left(-0.5Y_{it}^{0}+\eta_{it}\right)$, $\eta_{it}\sim N(0,0.5^{2})$.
@@ 4. DGP D:
    Treatment probability:
    $\Pr(D_{it}=1\mid Y_{it}^{0})=\mathrm{logit}\left(+0.5Y_{it}^{0}+\eta_{it}\right)$, $\eta_{it}\sim N(0,0.5^{2})$.
@@ 5. DGP E:
    Rare–extreme point mass (unseen support)
    $Y_{it}^{0}=\begin{cases}-10&\text{w.p.\ }0.002,\\ Z&\text{w.p.\ }0.996,\;Z\sim N(0,1),\\ +10&\text{w.p.\ }0.002,\end{cases}\qquad D_{it}\sim\mathrm{Bernoulli}(0.3).$
@@ 6. DGP F:
    $Y_{it}^{0}\sim\chi^{2}(3),\qquad D_{it}\sim\mathrm{Bernoulli}(0.3).$
@@ 7. DGP G:
DGP E is specifically constructed so that the finite sample has a high probability of not observing the population extrema, the precise scenario for which the hybrid band was developed.
@@ 5.1 Simulation Results
Table 5.1: Simultaneous coverage of the 95% hybrid and Manski bands
@@
A Standard normal 85.95 9.25 89.75 21.40 96.05 49.50
B $t_{3}$ heavy-tail 83.05 36.70 93.40 66.65 99.65 94.90
C AR(1) bias (–) 99.40 51.10 100 84.05 100 99.70
D AR(1) bias (+) 100 84.00 100 98.35 100 100
E Large extrema 89.10 27.30 93.65 46.45 98.80 81.80
F $\chi^{2}_{3}$ 100 99.85 100 100 100 100
G Uniform 100 100 100 100 100 100
Table 5.1 shows that the hybrid band fulfills its intended role whenever the finite sample is likely to miss the population extrema, while coinciding with the Manski interval in designs where the support is fully known. For the light-tailed Normal benchmark (DGP A) the plug-in Manski interval captures the true effect in barely one tenth of the B=2, [^0] B=2{,}000 replications for a sample size of $N=50$ and still under-covers at $N=100$. Adding the DKW pad and lifts hybrid coverage into the mid-80 percent range for $N=50$ and drives it close to the nominal 95 percent by $N=250$. The same pattern holds for the $t_{3}$ heavy-tail (DGP B): Manski improves as the sample begins to see extreme draws, but hybrid is uniformly closer to the target and reaches virtually perfect coverage in the largest sample.
Serial dependence and endogenous treatment (DGPs C and D) widen both bands. Under negative selection bias, Manski still misses the effect in half of the small-sample replications, whereas hybrid covers more than 99% of the time; with positive selection both bands converge, highlighting that the coverage gap arises specifically when the sample fails to capture the relevant tails. That point is most evident in the rare-extreme design (DGP E): the population includes outcomes of $\pm 10$ with probability only $0.2\%$, so the finite sample almost never observes them; Manski therefore covers only 27% at $N=50$, whereas the hybrid correction restores coverage to 89% and rises above 98% by $N=250$.
When the lower support is known to be zero, as under the left-truncated $\chi^{2}(3)$ baseline (DGP F), only the upper-tail pad is required and Manski already attains nominal coverage; hybrid is effectively the same band. The same coincidence is observed for the uniform distribution with fully known support (DGP G), where both methods hit 100 percent in every cell. Taken together, the results corroborate the theory: hybrid bands deliver the promised finite-sample protection precisely in situations where the classical plug-in Manski interval is too narrow and reduce to Manski when no tail uncertainty remains.
@@ 6 Empirical Application
In this section, we ask “Does gender-based board diversity causally affect firm innovation?”. We begin by outlining the data and summarizing its key descriptive statistics. We then present the nonparametric bounds approach of Manski, (1990, 2003) with simultaneous confidence bands in Proposition 1, the hybrid band proposed in Proposition 4, and the naïve mean‐comparison framework of Angrist and Pischke, (2009) (reported in Appendix D). The common objective is to test the null hypothesis of a zero average treatment effect of diversity on innovation—against a positive or negative alternative—when the diversity cut-off is selected endogenously (see Eq. (15)).
@@ 6.1 Data and Descriptive Statistics
In our analysis, we categorize the eleven GICS sectors into five broader groups: Cyclicals (Consumer Discretionary, Materials, Industrials, Real Estate), Defensives (Health Care, Consumer Staples, Utilities), Growth & Innovation (Information Technology, Communication Services), Financials, and Energy. This classification reflects the economic sensitivities of these sectors, as identified by MSCI. Specifically, MSCI’s Cyclical and Defensive Sectors Indexes classify sectors based on their performance correlation with the business cycle, using the OECD Composite Leading Indicator. According to MSCI, sectors like Consumer Discretionary, Materials, Industrials, Real Estate, Information Technology, Communication Services, and Financials are considered cyclical due to their positive correlation with economic expansions. Conversely, sectors such as Health Care, Consumer Staples, Utilities, and Energy are deemed defensive, exhibiting resilience during economic downturns. By adopting this grouping, we aim to capture the nuanced behaviors of these sectors in relation to macroeconomic conditions, facilitating a more informed analysis of sectoral dynamics. This classification can be found in table D.1.
Following the CSRD definition of a ‘large undertaking’ (Directive 2022/2464/EU, Art. 3 Pt 4) and the 250-employee threshold used in EU and UK gender-pay-gap statutes, we restrict the sample to firms whose time-average workforce is at least 250 employees over the sample horizon to ensure they fall under harmonized disclosure regimes. The restriction yields $n=901$ firms and a total of $N=25,228$ firm-quarter observations.
@@
Note: $N$ varies by variable because some firm–quarter observations are missing that particular item (e.g. Tobin’s $Q$ is reported for 23,085 of the 25,228 firm–quarters). The descriptive statistics are computed on all available values for each variable (“pair-wise” basis). For the causal analysis we use listwise deletion, retaining only the firm–quarters for which all diversity indicators and Tobin’s $Q$ are present.
@@
The key “treatment” variable is the percentage of women in senior leadership positions. These diversity measures are constructed using a supervised machine-learning algorithm applied to senior executives’ names, which infers gender from linguistic patterns. If the algorithm cannot assign a gender with high confidence, the individual is labeled as “unknown”. Importantly, the incidence of unknown classifications is very low: on average only about 0.03% (Table 6.1). The outcome of interest is Tobin’s $Q$, defined as the ratio of the firm’s market value to the replacement cost of its assets, a standard measure of firm performance and growth opportunities (Tobin,, 1969, 1978). We also utilize several control variables for descriptive analysis, including firm size (log total assets), leverage (debt-to-assets ratio), and total employees. Summary statistics for all main variables are provided in Table 6.1. After excluding observations with missing data on key fields, the average percentage of women in senior roles is about 27.5%. The standard deviation (around 12 percentage points for female share) indicates considerable cross-firm variation. Notably, a non-trivial subset of firm-quarters have zero diversity: roughly 4% of observations have no women in senior leadership, at least at some point in the sample. The distribution of the diversity variables is right-skewed. Figure 6.1 illustrates kernel density estimates of the percentage of female senior leaders across all firm-quarters. The distribution is skewed to the right with a primary mode around 25–35%, and a secondary mass at 0% corresponding to firms and periods with homogeneous leadership teams.
@@
Figure 6.1: Kernel density plot of percentage women. Scott’s rule (Scott,, 2015) is used to select the smoothing bandwidth parameter.
@@
We next explore the raw association between gender diversity measure and firm performance. In the full sample (pooled across all sectors and time periods), there is a strong positive correlation between senior-team diversity and Tobin’s $Q$. Figure 6.2 plots rolling correlations over time, using a moving window of half the sample period ($T/2\approx 14$ quarters) to track how the relationship evolves. The Pearson correlation between the percentage of women in leadership and Tobin’s $Q$ is in the range of +0.6 to +0.7 for most of the sample, indicating a fairly strong linear association. The figure also reports Kendall’s $\tau$ rank correlation, which captures monotonic association; this measure corroborates the positive link while being slightly lower in magnitude, suggesting the relationship is broadly monotonic even if not perfectly linear. The association appears to strengthen from 2015 up to about 2019, consistent with increasing awareness and implementation of diversity initiatives, but then shows a noticeable drop around 2019–2020. After 2019, the rolling correlations decline, implying that the previously tight diversity–performance relationship loosened and is increasing again after 2021. One possible interpretation is that external shocks or changing market conditions (for instance, the disruptive impact of the COVID-19 pandemic or the murder of George Floyd) temporarily weakened the correlation between diversity and market valuations.

### v2

@@ 4.1 Concentration-Driven Confidence Bands for Average Treatment Effects
  • The Dvoretzky-Kiefer-Wolfowitz concentration bound (which requires no tail assumptions beyond finiteness) for the order-statistic endpoints, and
  • The usual asymptotic delta‐method (CLT) for the sample means and proportions. The DKW inequality controls the uniform deviation $\sup_{x}\left\lvert F_{n}(x)-F(x)\right\rvert$ in finite samples without any distributional assumptions on $Y$ (see, e.g., Chapter 3 of van1996weak). This hybrid approach preserves the simplicity of the DKW envelope for the nonparametric piece while relying on asymptotic normality only for the low‐dimensional parameters.
Proposition 4 ($100(1-\alpha)\%$ hybrid confidence band under $\upalpha$-mixing).
Let $(Y_{ijt},Z_{ijt})_{i,j,t}$ be strictly stationary with $\upalpha$-mixing coefficients $\upalpha(r)$ such that
@@
$\varepsilon_{k}:=(1+4C_{\upalpha})\,\sqrt{\frac{2\log(8/\alpha_{u})}{N_{k}}},$
choose any $r_{k}\in\left(0,\tfrac{1}{2}-\varepsilon_{k}\right]$, and set the DKW-calibrated empirical endpoints
$L_{\alpha_{u}}^{(k)}:=Y^{(k)}_{\left(\lceil(r_{k}+\varepsilon_{k})N_{k}\rceil\right)},\quad U_{\alpha_{u}}^{(k)}:=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)},\qquad k=0,1,$
@@
$\sqrt{N}(\hat{\theta}-\theta)\xrightarrow{d}N(0,\Omega_{\theta}),$
for some positive semidefinite $4\times 4$ matrix $\Omega_{\theta}$. Let $\hat{\Omega}_{\theta}$ be any consistent estimator of $\Omega_{\theta}$ (e.g. a HAC or cluster-robust estimator).
Then a $100(1-\alpha)\%$ confidence band for the union of the bounds is
@@
where
$\nabla\mathcal{L}(\theta)$ $=\left(p_{1},\;-p_{0},\;\delta_{1}-U_{\alpha_{u}}^{(0)},\;L_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top},$
$\nabla\mathcal{U}(\theta)$ $=\left(p_{1},\;-p_{0},\;\delta_{1}-L_{\alpha_{u}}^{(0)},\;U_{\alpha_{u}}^{(1)}-\delta_{0}\right)^{\top},$
and
$\text{S.E.}\left(\mathcal{L}_{\alpha_{u}}(\hat{\theta})\right):=\sqrt{\frac{1}{N}\nabla\mathcal{L}(\hat{\theta})^{\top}\hat{\Omega}_{\theta}\nabla\mathcal{L}(\hat{\theta})},\quad\text{S.E.}\!\left(\mathcal{U}_{\alpha_{u}}(\hat{\theta})\right):=\sqrt{\frac{1}{N}\nabla\mathcal{U}(\hat{\theta})^{\top}\hat{\Omega}_{\theta}\nabla\mathcal{U}(\hat{\theta})}$.
Proposition 4 thus guarantees that our hybrid concATE band achieves the desired confidence level for the ATE bounds even when data are dependent, by using the DKW inequality with appropriate mixing corrections.
Following the same logic as Corollary 1, the extension of Proposition 4 to the case of one-tail truncated latent conditional expectations replaces the endpoints by
$L^{(k)}:=\lambda,\qquad U^{(k)}_{\alpha_{u}}:=\hat{F}^{(k),-1}_{N_{k}}\left(1-r_{k}-\tilde{\varepsilon}_{k}\right)=Y^{(k)}_{\left(\lceil(1-r_{k}-\tilde{\varepsilon}_{k})N_{k}\rceil\right)},$
where $\tilde{\varepsilon}_{k}=(1+4C_{\upalpha})\sqrt{2\log(4/\alpha_{u})/N_{k}}$ and $r_{k}\in(0,1-\tilde{\varepsilon}_{k}]$.
@@ 5.1 Monte Carlo Design
correspondingly for the Manski and Hybrid approaches.
Within each replication and arm $k\in\{0,1\}$ we compute
$\varepsilon_{k}:=\sqrt{\frac{\log\left(8/\alpha\right)}{2N_{k}}},\quad r_{k}\in\left(0,\frac{1}{2}-\varepsilon_{k}\right],$
and set the outward empirical-quantile endpoints
$L^{(k)}:=Y^{(k)}_{\left(\lceil(r_{k}+\varepsilon_{k})N_{k}\rceil\right)},\quad U^{(k)}:=Y^{(k)}_{\left(\lceil(1-r_{k}-\varepsilon_{k})N_{k}\rceil\right)}.$
(When $\varepsilon_{k}\leq\frac{1}{4}$ we take $r_{k}=\varepsilon_{k}$, yielding the convenient indices $\lceil 2\varepsilon_{k}N_{k}\rceil$ and $\lceil(1-2\varepsilon_{k})N_{k}\rceil$; if $\varepsilon_{k}>\frac{1}{4}$, we set $r_{k}=\frac{1}{2}-\varepsilon_{k}-\frac{1}{N_{k}}$ to preserve ordering.) The Manski benchmark uses the usual plug-in with known support when available (DGP G), and otherwise the sample extrema. The realised outcome is
@@ 3. DGP C:
    Treatment probability: $\Pr(D_{it}=1\mid Y_{it}^{0})=\mathrm{logit}\!\left(-0.5Y_{it}^{0}+\eta_{it}\right)$, $\eta_{it}\sim N(0,0.5^{2})$.
@@ 4. DGP D:
    Treatment probability: $\Pr(D_{it}=1\mid Y_{it}^{0})=\mathrm{logit}\!\left(+0.5Y_{it}^{0}+\eta_{it}\right)$, $\eta_{it}\sim N(0,0.5^{2})$.
@@ 5. DGP E:
    Rare-extreme point mass (controlled tail visibility)
    Let the per-tail probability be $\pi_{N}=\lambda/(2N)$ with $\lambda\in\{0.35,\,0.69,\,1.61\}$ so that $\Pr(\text{no extreme in the sample})\approx e^{-\lambda}\in\{0.70,\,0.50,\,0.20\}$. Set
    $Y_{it}^{0}=\begin{cases}-M,&\text{w.p.\ }\pi_{N},\\ Z,&\text{w.p.\ }1-2\pi_{N},\quad Z\sim N(0,1),\\ +M,&\text{w.p.\ }\pi_{N},\end{cases}\quad M\in\{6,10\},\quad D_{it}\sim\mathrm{Bernoulli}(0.3).$
    Purpose: directly tunes the probability that finite samples miss the true extremes.
@@ 6. DGP F:
    $Y_{it}^{0}\sim\chi^{2}(3)$, $D_{it}\sim\mathrm{Bernoulli}(0.3)$.
    Note: for the hybrid band we fix $L^{(k)}=0$ and use only the one-sided upper empirical quantile $U^{(k)}=\hat{F}^{-1}_{N_{k}}\left(1-r_{k}-\varepsilon_{k}\right)$.
@@ 7. DGP G:
DGP E is specifically constructed so that the finite sample has a tunable probability of not observing the population extremes, the precise scenario for which the hybrid band was developed.
@@ 5.2 Simulation Results
Table 5.1: Pointwise Monte-Carlo coverage of the 95% hybrid and Manski bounds at $\tau^{\circ}=50\%$
@@
A Standard normal 99.20 74.9 99.80 79.40 100 93.80
B $t_{3}$ heavy-tail 99.90 75.50 100 90 100 99.50
C AR(1) bias (-) 92.50 98.30 83.80 99.70 100 100
D AR(1) bias (+) 99.80 99.80 96.50 100 100 100
E Large extrema (controlled) 99.50 77.80 99.60 87.70 100 97.70
F $\chi^{2}_{3}$ (left-trunc.) 99.50 100 99.80 100 100 100
G Uniform (known support) 100 100 100 100 100 100
Table 5.1 reports the pointwise Monte-Carlo coverage of the identified set for the treatment effect evaluated at the tipping point $\tau^{\circ}=50\%$. For each replication, we check whether $\Delta$ lies inside the Manski and the Hybrid bounds at this single threshold. This is not a simultaneous functional coverage result; it is pointwise identification coverage at $\tau^{\circ}$.
For the light-tailed Normal benchmark (DGP A) the plug-in Manski interval under-covers at small and moderate $N$ (75–79% at $N\in\{50,100\}$, 94% at $N=250$), whereas replacing sample extrema by DKW-calibrated outward empirical quantiles yields coverage essentially at nominal across all $N$ (99–100%). The same pattern holds for the $t_{3}$ heavy tail (DGP B): Manski improves with $N$ (76%, 90%, 99.5%), while the hybrid band remains near 100% throughout.
Serial dependence and endogenous treatment (DGPs C and D) widen both bands. Under negative selection bias (DGP C) the hybrid under-covers at smaller $N$ (93% at $N=50$, 84% at $N=100$) relative to Manski (98–100%), converging by $N=250$; this is consistent with using the i.i.d. DKW radius under dependence and disappears with the mixing-adjusted calibration. With positive selection (DGP D) both bands are near-nominal and coincide by $N=250$.
The advantage of the hybrid is most evident in the rare-extreme design (DGP E): small and moderate samples frequently miss the population extremes, so Manski under-covers (78–88%), whereas the quantile-calibrated hybrid correction restores coverage to about 100% across all $N$.
When the lower support is known to be zero, as under the left-truncated $\chi^{2}(3)$ baseline (DGP F), only the upper empirical quantile is required and both methods are effectively the same (99.5–100%). The same coincidence is observed for the uniform distribution with fully known support (DGP G), where both methods hit 100% in every cell. Taken together, the results corroborate the theory: hybrid bands deliver the promised finite-sample protection precisely in situations where the classical plug-in Manski interval is too narrow, and they reduce to Manski when no tail uncertainty remains.
@@ 6 Empirical Application
In this section, we ask “Does gender-based board diversity causally affect firm innovation?”. We begin by outlining the data and summarizing its key descriptive statistics. We then present the nonparametric bounds approach of manski1990nonparametric; manski2003partial with simultaneous confidence bands in Proposition 1, the hybrid band proposed in Proposition 4, and the naïve mean‐comparison framework of angrist2009mostly (reported in Appendix D). The common objective is to test the null hypothesis of a zero average treatment effect of diversity on innovation, against a positive or negative alternative, when the diversity cut-off is selected endogenously (see Eq. (15)).
@@ 6.1 Data and Descriptive Statistics
In our analysis, we categorize the eleven Global Industry Classification Standard sectors (GICS hereafter) into five broader groups: Cyclicals (Consumer Discretionary, Materials, Industrials, Real Estate), Defensives (Health Care, Consumer Staples, Utilities), Growth & Innovation (Information Technology, Communication Services), Financials, and Energy. This classification reflects the economic sensitivities of these sectors, as identified by Morgan Stanley Capital International (MSCI hereafter). Specifically, MSCI’s Cyclical and Defensive Sectors Indexes classify sectors based on their performance correlation with the business cycle, using the OECD Composite Leading Indicator. According to MSCI, sectors like Consumer Discretionary, Materials, Industrials, Real Estate, Information Technology, Communication Services, and Financials are considered cyclical due to their positive correlation with economic expansions. Conversely, sectors such as Health Care, Consumer Staples, Utilities, and Energy are deemed defensive, exhibiting resilience during economic downturns. By adopting this grouping, we aim to capture the nuanced behaviours of these sectors in relation to macroeconomic conditions, facilitating a more informed analysis of sectoral dynamics. This classification can be found in Table D.1.
Following the Corporate Sustainability Reporting Directive definition (CSRD hereon) of a ‘large undertaking’ (Directive 2022/2464/EU, Art. 3 Pt 4) and the 250-employee threshold used in EU and UK gender-pay-gap statutes, we restrict the sample to firms whose time-average workforce is at least 250 employees over the sample horizon to ensure they fall under harmonized disclosure regimes. The restriction yields $n=901$ firms and a total of $N=25,228$ firm-quarter observations.
@@
Note: $N$ varies by variable because some firm-quarter observations are missing that particular item (e.g. Tobin’s $Q$ is reported for 23,085 of the 25,228 firm-quarters). The descriptive statistics are computed on all available values for each variable (“pair-wise” basis). For the causal analysis we use list-wise deletion, retaining only the firm-quarters for which all diversity indicators and Tobin’s $Q$ are present. Tobin’s $Q$ is scaled using a robust scaler (median and interquartile range) prior to the causal analysis.
@@
The key “treatment” variable is the percentage of women in senior leadership positions. These diversity measures are constructed using a supervised machine-learning algorithm applied to senior executives’ names, which infers gender from linguistic patterns. If the algorithm cannot assign a gender with high confidence, the individual is labelled as “unknown”. Importantly, the incidence of unknown classifications is very low: on average only about 0.03% (Table 6.1). The outcome of interest is Tobin’s $Q$, defined as the ratio of the firm’s market value to the replacement cost of its assets, a standard measure of firm performance and growth opportunities (tobin1969general; tobin1978monetary). We also utilize several control variables for descriptive analysis, including firm size (log total assets), leverage (debt-to-assets ratio), and total employees. Summary statistics for all main variables are provided in Table 6.1. After excluding observations with missing data on key fields, the average percentage of women in senior roles is about 27.5%. The standard deviation (around 12 percentage points for female share) indicates considerable cross-firm variation. Notably, a non-trivial subset of firm-quarters have zero diversity: roughly 4% of observations have no women in senior leadership, at least at some point in the sample. The distribution of the diversity variables is right-skewed. Figure 6.1 illustrates kernel density estimates of the percentage of female senior leaders across all firm-quarters. The distribution is skewed to the right with a primary mode around 25–35%, and a secondary mass at 0% corresponding to firms and periods with homogeneous leadership teams.
@@
Figure 6.1: Kernel density plot of percentage women. Scott’s rule (scott2015multivariate) is used to select the smoothing bandwidth parameter.
@@
We next explore the raw association between gender diversity measure and firm performance. In the full sample (pooled across all sectors and time periods), there is a strong positive correlation between senior-team diversity and Tobin’s $Q$. Figure 6.2 plots rolling correlations over time, using a moving window of half the sample period ($T/2\approx 14$ quarters) to track how the relationship evolves. The Pearson correlation between the percentage of women in leadership and Tobin’s $Q$ is in the range of +0.6 to +0.7 for most of the sample, indicating a fairly strong linear association. The Figure also reports Kendall’s $\tau$ rank correlation, which captures monotonic association; this measure corroborates the positive link while being slightly lower in magnitude, suggesting the relationship is broadly monotonic even if not perfectly linear. The association appears to strengthen from 2015 up to about 2019, consistent with increasing awareness and implementation of diversity initiatives, but then shows a noticeable drop around 2019–2020. After 2019, the rolling correlations decline, implying that the previously tight diversity-performance relationship loosened and is increasing again after 2021. One possible interpretation is that external shocks or changing market conditions (for instance, the disruptive impact of the COVID-19 pandemic or the murder of George Floyd) temporarily weakened the correlation between diversity and market valuations.